# Parastatistics revealed: Peierls phase twists and shifted conformal towers in interacting periodic chains

### v1

@@ Abstract
We consider interacting paraparticle chains with a constant $R$-matrix where the Hamiltonian sums over the internal degrees (flavors) of the paraparticles. For such flavor-blind Hamiltonians we show a general factorization of the Hilbert space into occupation and flavor parts with the Hamiltonian acting non-trivially only on the former. For open boundaries, the spectrum therefore coincides with that of the occupation Hamiltonian $H_{\rm occ}$ with the flavor part merely adding degeneracies. For periodic boundaries, a cyclic reordering of the flavors leads to a separation of $H_{\rm occ}$ into flux sectors at fixed particle number, thus making the parastatistics directly observable in the energy spectrum. For important exemplary cases, $H_{\rm occ}$ reduces to the XXZ chain with flux allowing for an exact solution. In the gapless regime, this solution shows flux-shifted $c=1$ conformal towers in the low-energy spectrum and a temperature-dependent chemical potential in the bulk thermodynamics.
@@
Introduction—When considering identical quantum particles, the distinction between bosons and fermions is most fundamental. For example, it manifests itself in the formation of Bose–Einstein condensates in ultracold atomic gases and Fermi surfaces in solids, which both in turn dictate the macroscopic physical properties. At a more technical level, bosons and fermions are distinguished by the behavior of the many-body wave functions under particle permutations, which can be encoded in the commutation relations of the respective creation and annihilation operators [1].
@@
Given the profound consequences of different quantum statistics, there have been various efforts to go beyond bosons and fermions. To give just a few examples, in two dimensions one can consider braiding instead of permutations [2], leading to so-called anyons [3, 4] which emerge as quasiparticles in fractional quantum Hall states [5, 6, 7]. Anyonic statistics is also displayed by $\mathbb{Z}_{p}$-parafermions [8, 9]—generalizations of Majorana fermions in Potts/clock models—which are suspected to exist in quantum Hall/superconductor hybrid systems and topological insulators [10, 11]. Furthermore, building on parafermions one can construct so-called Fock parafermions [12, 13, 14], which show a generalized Pauli principle in the sense that single-particle levels can be occupied by at most $p$ particles. In fact, exclusion statistics even with fractional statistical parameters can be defined in any dimension [15, 16] and is realized by spinons in antiferromagnetic spin chains [17] and conformal field theories [18]. However, all generalizations mentioned above appear only in strongly correlated systems and thus are not amenable to simple single-particle descriptions. A notable exception is the Baxter–Fendley model [19, 20], whose many-particle spectrum is built from single-particle energies of $\mathbb{Z}_{p}$-parafermions. However, the model is non-Hermitian, thus obscuring its quantum mechanical interpretation.
@@
Very recently, Wang and Hazzard [21] introduced a parastatistics based on non-trivial bilinear relations between the second quantized operators acting on internal flavors, in contrast to Green-type trilinear algebras [23, 24], which complicate thermodynamic derivations [25]. For the newly introduced paraparticles, Wang and Hazzard studied the generalized exclusion and exchange statistics and derived the exact energy spectra of certain flavor-blind bilinear Hamiltonians by relating them to specific quantum spin chains with open boundary conditions (OBC). In this sense, for flavor-blind bilinear chains with OBC, the Hamiltonians realize models of free paraparticles whose many-body spectra are built from single-particle energies with a generalized exclusion principle encoded in the mode multiplicities $d_{n}$ [21, 26].
@@
In this Letter we go beyond free paraparticle chains with OBC by considering chains with periodic boundary conditions (PBC) and interactions. In contrast to the OBC case, already the single-particle energies are modified by a Peierls twist due to the parastatistics for periodic boundaries. Our main general results are: (i) a generic factorization theorem for the Hilbert space of flavor-blind Hamiltonians, (ii) an explicit derivation of the differences between OBC and PBC: with OBC the fixed flavor order gives only degeneracies while with PBC the cyclic permutation of flavors induces a Peierls twist in $H_{\rm occ}$, splitting the energy spectrum into flux sectors, and (iii) an exact formula for the dimension of the flavor subspaces in terms of character projectors of the cyclic group. For the specific case that the single-mode occupations are given by $d_{0}=1$, $d_{1}=m$, and $d_{n}=0$ for $n\geq 2$ we show, furthermore: (iv) that the occupation sector for paraparticles with nearest-neighbor interaction is the Bethe-ansatz-solvable XXZ chain with flux. This exact solution shows (v) that for PBC the low-energy spectrum consists of conformal towers with a shift between persistent current branches, and (vi) that the bulk thermodynamics shows two signatures of the paraparticle character of the constituent particles: a zero-temperature residual entropy and a temperature-dependent chemical potential.
@@
General setup—We use the second quantized formulation of parastatistics recently introduced in Ref. [21]. Specifically we consider a one-dimensional chain of $L$ lattice sites, at each of which we define operators $\psi_{i,a}^{\pm}$, $i=1,\ldots,L$, that create or annihilate a particle with internal flavor $a=1,\ldots,F$. The parastatistics is encoded in the commutation relations
$\hskip-11.38109pt\psi_{i,a}^{+}\psi_{j,b}^{+}=\sum_{cd}R^{cd}_{ab}\psi_{j,c}^{+}\psi_{i,d}^{+},\quad\psi_{i,a}^{-}\psi_{j,b}^{-}=\sum_{cd}R^{ba}_{dc}\psi_{j,c}^{-}\psi_{i,d}^{-},$
$\hskip-11.38109pt\psi^{-}_{i,a}\psi^{+}_{j,b}=\sum_{cd}R^{ac}_{bd}\psi^{+}_{j,c}\psi^{-}_{i,d}+\delta_{ab}\delta_{ij},$ (1)
where $R_{ab}^{cd}$ are $F\times F$ matrices satisfying $\sum_{\sigma\tau}R_{ab}^{\sigma\tau}R_{\sigma\tau}^{cd}=\delta_{a}^{c}\delta_{b}^{d}$ and $\sum_{\sigma\tau\kappa}R_{ab}^{\sigma\tau}R_{\tau c}^{\kappa u}R_{\sigma\kappa}^{de}=\sum_{\sigma\tau\kappa}R_{bc}^{\sigma\tau}R_{a\sigma}^{d\kappa}R_{\kappa\tau}^{eu}$. (We note that the $R$-matrix defined in (1) is known as the permuted $R$-matrix in the literature of integrable systems [28], with the latter relation being the constant Yang–Baxter equation.) In the special case $R_{ab}^{cd}=\pm\delta_{a}^{d}\delta_{b}^{c}$ the relations (1) simplify to bosons and fermions with $F$ internal degrees of freedom.
Using these second quantized operators we introduce the OBC or PBC Hamiltonian
@@
where $n_{i}=\sum_{a}n_{i,a}$ and $n_{i,a}=\psi^{+}_{i,a}\psi^{-}_{i,a}$ are the total and flavor particle density, respectively. We note that in analogy to Ref. [21] a sum over the internal flavor degrees is performed, making the Hamiltonian flavor-blind. In contrast to Ref. [21], we do include an explicit interaction term which is allowed to be long range for the general discussion below. In the specific examples considered at the end of this letter, the interaction will be limited to nearest-neighbors only. For a single mode, we denote the number of possible states with $n$ particles by $d_{n}$ [22].
Hilbert space structure—The first important point to realize is that for flavor-blind Hamiltonians such as Eq. (2) the Hilbert space factorizes.
Factorization theorem: We can always write the local Hilbert space at site $i$ as
$\mathcal{H}_{i}=\bigoplus_{n=0}^{n_{\rm max}}|n_{i}\rangle\otimes\mathcal{F}_{n}$ (3)
with $\mbox{dim}\,\mathcal{F}_{n}=d_{n}$. The multiplicities $d_{n}$ depend on the considered model and are, in general, not equal to the algebraic dimension $F$ of the internal vector space of the $R$ matrices. The Hamiltonian only acts on the occupation number part $\{|n_{i}\rangle\}$ while the flavor part is a ’bystander’ creating only additional degeneracies in the many-body spectrum. For OBC the order of the flavors cannot change and the degeneracy is simply the dimension of the flavor space compatible with the occupation number eigenstate. A general state at fixed particle number $N$ is given by $|\Psi\rangle_{N}=\sum_{s}a_{s}|\{n_{i}\}_{s}\rangle\otimes|\phi_{\rm fl}\rangle$ with $\sum_{i}n_{i}=N$ for each $s$. Each occupation number configuration $|\{n_{i}\}\rangle$ has a compatible flavor space $\mathcal{F}(\{n_{i}\})=\bigotimes_{i}\mathcal{F}_{n_{i}}$.
OBC degeneracy rule: The dimension of the flavor space belonging to the state $|\Psi\rangle_{N}$ is $D=\mbox{dim}\left[\bigcap_{a_{s}\neq 0}\mathcal{F}(\{n_{i}\}_{s})\right]$.
@@
This all changes when we consider PBC. In this case we can wrap a flavor around the chain thus obtaining cyclic permutations in the order of the flavors. To understand how the eigenstates of the paraparticle Hamiltonian can be constructed from an occupation number Hamiltonian plus flavor degeneracies, one then has to study some basic properties of the cyclic group first. We define $N_{n}$ as the number of particles occupying modes with degeneracy $d_{n}$. Then $N=\sum_{n}N_{n}=\sum_{i}n_{i}$ is the total particle number. We define, furthermore, $M=M(\{N_{n}\})$ as the total number of flavors which actually does get cyclically permuted under PBC. It is important to note that the length of the flavor string $M$ is in general not equal to the number of particles $N$. Consider, for example, the case $d_{0}=1$, $d_{1}=2$, $d_{2}=1$, and $d_{n}=0$ for $n>2$. In this case $M=N_{1}$, where $N_{1}$ is the number of singly occupied modes because the vacuum and the doubly occupied modes do not have a flavor index.
@@
with $\delta\equiv\gamma_{q}$. I.e., the occupation number wavefunction $|\Psi_{\rm occ}\rangle$ picks up a phase which is equal and opposite to the phase picked up by the flavor part. The eigenspaces of the occupation number Hamiltonian for PBC thus separate into spaces with Peierls phases $\gamma_{q}$ which is equivalent to saying that there is now a flux penetrating the ring. As a consequence, the paraparticle statistics—which is responsible for the phases $\gamma_{q}$—is directly reflected in the eigenspace structure and energies of the occupation number Hamiltonian $H_{\rm occ}$ for PBC.
@@
Example—To demonstrate these general results, we consider from now on example 3 of Ref. [21]. In this hardcore case there is only one type of flavor which is associated with the single occupancy of a mode and every particle carries one of the $m$ possible flavor labels implying $M=M_{1}=N$. In addition to the hopping terms and chemical potential considered in the reference above, we also allow for a nearest-neighbor density-density interaction with strength $V$. The paraparticle chain can then be mapped onto a spin chain
@@
with $n_{i}=\sum_{\sigma}S^{+}_{i\sigma}S^{-}_{i\sigma}$ where $i$ denotes the sites of the lattice and $\sigma=1,\cdots,m$ is the flavor index. Note that in order to obtain an exact solution we have chosen the hopping $J$, chemical potential $\mu$, and interaction $V$ to be site independent. The following separation into occupation number and flavor part, however, also holds for site-dependent parameters. For this particular model, the separation into these two parts is achieved easily by embedding the Hilbert space into a larger one, $\mathcal{H}\subset\mathcal{H}^{\rm occ}\otimes\mathcal{H}^{\rm fl}$. Here the local occupation Hilbert space is two-dimensional $H^{\rm occ}_{i}=\mbox{span}(|0\rangle,|1\rangle)$ while the local flavor space is $m$-dimensional, $\mathcal{H}_{i}^{\rm fl}=\mbox{span}(|s_{1}\rangle,\cdots,|s_{m}\rangle)$. To achieve the embedding we simply identify $|0\rangle\equiv|0,s_{1}\rangle$. The Hamiltonian then becomes
@@
with $N$ being the particle number. In a block with $N$ fixed, the chemical potential only contributes a constant but does not affect the eigenvectors. The last term is only present for PBC in which case the Peierls phase is $\gamma_{q}(N)=2\pi q/N$ which can also be distributed uniformly between all bonds. For OBC there is thus no separation into different flux sectors and the dimension of the flavor sector for fixed NN—which determines the degeneracy of each XXZ eigenvalue—is simply $D=m^{N}$. For PBC, we can use the general degeneracy formula (Parastatistics revealed: Peierls phase twists and shifted conformal towers in interacting periodic chains) with $c_{1}=g=\mbox{gcd}(N,r)$ and $d_{1}=m$ leading to
@@
While the many-body eigenenergies in the interacting case can be obtained by the Bethe ansatz—or by numerical methods for cases where the occupation number Hamiltonian is not integrable—they can be constructed in the non-interacting case, $V=0$, from the single particle eigenenergies alone. For OBC, these energies are $\varepsilon_{k}=-2J\cos k_{r}-\mu$ with $k_{r}=\frac{\pi r}{L+1}$, and $r=1,\cdots,L$. The many-body energies are then given by $E(\{n_{k}\})=\sum_{k}n_{k}\varepsilon_{k}$ with $n_{k}\in\{0,1\}$, $\sum_{k}n_{k}=N$, $N=1,\cdots,L$ and each energy with $N$ particles has flavor degeneracy $m^{N}$. For PBC, on the other hand, the single-particle eigenvalues are $\varepsilon_{k}(N,q)=-2J\cos k_{r,q}-\mu$ with $k_{r,q}=\frac{2\pi r+\gamma_{q}}{L}$, $r=0,\cdots,L-1$, and $q=0,\cdots,N-1$. The many-body eigenstates are obtained by going through all allowed values of $N,q$ and in each case constructing all possible eigenvalues $E(\{n_{k}\})=\sum_{k}n_{k}\varepsilon_{k}(N,q)$ with $\sum_{k}n_{k}=N$ and with each one of them having an additional degeneracy of $\mbox{dim}\,\mathcal{H}^{\rm fl}_{N,q}$, see Eq. (9).
Bosonization—For the flavor-blind Hamiltonians (2) investigated here, the non-trivial part of the low-energy physics is entirely determined by the occupation number part of the Hamiltonian. This part can often be described by ordinary fermions, making it possible to classify the universal behavior using standard techniques. To be concrete, we continue with our example where the occupation number part of the Hamiltonian with PBC is an XXZ chain with a Peierls phase. For $|V/J|<2$, every sector with $0<N<L$ will be gapless and described by a conformal field theory with central charge $c=1$.
@@
Here $e_{\infty}$ is the energy per site in the thermodynamic limit, $L\to\infty$. The first term on the r.h.s. is the well known universal finite-size correction with central charge $c=1$ and $v$ the velocity of the excitations [33, 34]. The second term are the oscillator modes with $N_{n}^{R/L}$ counting the occupation of the $n$-th mode for right/left-movers, respectively. The parastatistics enters through the third term. In the low-energy effective theory, the Peierls phase $\gamma_{q}(N)=2\pi q/N$ leads to a persistent current $I(q)=-\frac{\pi vK}{L}(J-q/N)$ with $J=0$ if $q<\lfloor N/2\rfloor$ and $J=1$ if $q>\lfloor N/2\rfloor$. If $q=N/2$ then we have a degeneracy corresponding to equal and opposite persistent currents. In the finite-size spectrum the presence of these persistent currents means that we have conformal towers shifted by $\Delta E=\frac{2\pi vK}{L}\min\left(\frac{q}{N},1-\frac{q}{N}\right)^{2}$ (for $q$ and NqN-q the shift is the same, corresponding to opposite persistent currents) providing a clear signature of the parastatistics in the energy spectrum for PBC. Each of these levels will carry an additional degeneracy of $\mbox{dim}\,\mathcal{H}^{\rm fl}_{N,q}$ due to the flavor sector, see Eq. (9). The parameters $v$ and $K$ (with $K=1$ in the free case) are known exactly from the Bethe ansatz solution of the XXZ model for arbitrary filling.
Thermodynamics—In the thermodynamic limit, the $1/L$-terms due to the boundary conditions become irrelevant. This means that we can always start from the open boundary case when considering the $L\to\infty$ limit where the parastatistics only leads to additional degeneracies due to the flavor sector while the occupation number sector is not affected. Returning to our example, this means that the free energy per site is given by ($\beta=1/T$)

### v3

# arXiv:2511.01150v3[cond-mat.stat-mech] 28 Feb 2026

@@ Abstract
We consider interacting paraparticle chains with a constant $R$-matrix where the Hamiltonian sums over the internal degrees (flavors) of the paraparticles. For such flavor-blind Hamiltonians, we show a general factorization of the Hilbert space into occupation and flavor parts with the Hamiltonian acting non-trivially only on the former. For open boundaries, the spectrum therefore coincides with that of the occupation Hamiltonian $H_{\rm occ}$ with the flavor part merely adding degeneracies. For periodic boundaries, a cyclic reordering of the flavors leads to a separation of $H_{\rm occ}$ into flux sectors at fixed particle number, thus making the parastatistics directly observable in the energy spectrum. For important exemplary cases, $H_{\rm occ}$ reduces to the XXZ chain with flux, allowing for an exact solution. In the gapless regime, this solution shows flux-shifted $c=1$ conformal towers in the low-energy spectrum and a temperature-dependent chemical potential in the bulk thermodynamics.
@@
Introduction—When considering identical quantum particles, the distinction between bosons and fermions is of fundamental importance. For example, it manifests itself in the formation of Bose–Einstein condensates in ultracold atomic gases and Fermi surfaces in solids, both of which in turn dictate the macroscopic physical properties. At a more technical level, bosons and fermions are distinguished by the behavior of the many-body wave functions under particle permutations, which can be encoded in the commutation relations of the respective creation and annihilation operators [1].
@@
Given the profound consequences of different quantum statistics, there have been various efforts to go beyond bosons and fermions. To give just a few examples, in two dimensions one can consider braiding instead of permutations [2], leading to so-called anyons [3, 4], which emerge as quasiparticles in fractional quantum Hall states [5, 6, 7]. Anyonic statistics is also displayed by $\mathbb{Z}_{p}$-parafermions [8, 9]—generalizations of Majorana fermions in Potts/clock models—which are suspected to exist in quantum Hall/superconductor hybrid systems and topological insulators [10, 11]. Furthermore, building on parafermions one can construct so-called Fock parafermions [12, 13, 14], which show a generalized Pauli principle in the sense that single-particle levels can be occupied by at most $p$ particles. In fact, exclusion statistics even with fractional statistical parameters can be defined in any dimension [15, 16] and is realized by spinons in antiferromagnetic spin chains [17] and conformal field theories [18]. However, all generalizations mentioned above appear only in strongly correlated systems and thus are not amenable to simple single-particle descriptions. A notable exception is the Baxter–Fendley model [19, 20], whose many-particle spectrum is built from single-particle energies of $\mathbb{Z}_{p}$-parafermions. However, the model is non-Hermitian, thus obscuring its quantum mechanical interpretation.
@@
Very recently, Wang and Hazzard [21] introduced a parastatistics based on non-trivial bilinear relations between the second-quantized operators acting on internal flavors, in contrast to Green-type trilinear algebras [23, 24], which complicate thermodynamic derivations [25]. For the newly introduced paraparticles, Wang and Hazzard studied the generalized exclusion and exchange statistics and derived the exact energy spectra of certain flavor-blind bilinear Hamiltonians by relating them to specific quantum spin chains with open boundary conditions (OBC). In this sense, for flavor-blind bilinear chains with OBC, the Hamiltonians realize models of free paraparticles whose many-body spectra are built from single-particle energies with a generalized exclusion principle encoded in the mode multiplicities $d_{n}$ [21, 26].
@@
In this Letter, we go beyond free paraparticle chains with OBC by providing the first systematic study of interacting paraparticle chains with periodic boundary conditions (PBC). In contrast to the OBC case, the single-particle energies are already modified by a Peierls twist due to the parastatistics for periodic boundaries. Importantly, this twist is not imposed externally but emerges solely from the statistics encoded in the quadratic R-algebra. Our main general results are: (i) a generic factorization theorem for the Hilbert space of flavor-blind Hamiltonians, (ii) an explicit derivation of the differences between OBC and PBC: with OBC the flavors only contribute degeneracies while with PBC the cyclic permutation of flavors induces a Peierls twist in $H_{\rm occ}$, splitting the energy spectrum into flux sectors, and (iii) an exact formula for the dimension of the flavor subspaces in terms of character projectors of the cyclic group. For the specific case that the single-mode occupations are given by $d_{0}=1$, $d_{1}=m$, and $d_{n}=0$ for $n\geq 2$, we show, furthermore (iv) that the occupation sector for paraparticles with nearest-neighbor interaction is the Bethe-ansatz-solvable XXZ chain with flux. This exact solution shows (v) that for PBC the low-energy spectrum consists of conformal towers with a shift between persistent current branches, and (vi) that the bulk thermodynamics shows two signatures of the paraparticle character of the constituent particles: a zero-temperature residual entropy and a temperature-dependent chemical potential. Together, these results establish interacting periodic chains as concrete benchmark systems in which R-parastatistics leads to directly observable many-body signatures.
@@
General setup—We use the second-quantized formulation of parastatistics recently introduced in Ref. [21]. Specifically, we consider a one-dimensional chain of $L$ lattice sites, at each of which we define operators $\psi_{i,a}^{\pm}$, $i=1,\ldots,L$, that create or annihilate a particle with internal flavor $a=1,\ldots,F$. The parastatistics is encoded in the commutation relations
$\hskip-11.38109pt\psi_{i,a}^{+}\psi_{j,b}^{+}=\sum_{cd}R^{cd}_{ab}\psi_{j,c}^{+}\psi_{i,d}^{+},\quad\psi_{i,a}^{-}\psi_{j,b}^{-}=\sum_{cd}R^{ba}_{dc}\psi_{j,c}^{-}\psi_{i,d}^{-},$ (1)
$\hskip-11.38109pt\psi^{-}_{i,a}\psi^{+}_{j,b}=\sum_{cd}R^{ac}_{bd}\psi^{+}_{j,c}\psi^{-}_{i,d}+\delta_{ab}\delta_{ij},$
where $R_{ab}^{cd}$ are $F^{2}\times F^{2}$ matrices satisfying $\sum_{\sigma\tau}R_{ab}^{\sigma\tau}R_{\sigma\tau}^{cd}=\delta_{a}^{c}\delta_{b}^{d}$ and $\sum_{\sigma\tau\kappa}R_{ab}^{\sigma\tau}R_{\tau c}^{\kappa u}R_{\sigma\kappa}^{de}=\sum_{\sigma\tau\kappa}R_{bc}^{\sigma\tau}R_{a\sigma}^{d\kappa}R_{\kappa\tau}^{eu}$. In the following we also assume unitarity, $\sum_{\sigma\tau}R_{ab}^{\sigma\tau}(R_{cd}^{\sigma\tau})^{*}=\delta_{ac}\delta_{bd}$, so that $(\psi_{i,a}^{+})^{\dagger}=\psi_{i,a}^{-}$ and $H$ is Hermitian with a real spectrum and a standard Gibbs ensemble. We note that the $R$-matrix defined in (1) is known as the permuted $R$-matrix in the literature of integrable systems [27], with the latter relation being the constant Yang–Baxter equation. In the special case $R_{ab}^{cd}=\pm\delta_{a}^{d}\delta_{b}^{c}$ the relations (1) simplify to bosons and fermions with $F$ internal degrees of freedom. The constant YBE above defines braid generators $B_{i}$ built from $R$ acting on neighboring flavor spaces and satisfying the braid relations. The first quadratic condition below Eq. (1) implies $R^{2}=1$, so that $B_{i}^{2}=1$ and the braid representation reduces to a representation of the symmetric group. Exchange eigenvalues are therefore restricted to $\pm 1$, and no braid-type exchange statistics can arise within this class. The allowed statistics thus consists of bosonic and fermionic exchange channels which can be coupled by local constraints, leading to generalized exclusion rules and flavor-dependent degeneracies beyond simple independent bosons and fermions (see End Matter). For $F=2$, all constant solutions of the YBE have been classified by Hietarinta [28], and we analyze the corresponding generalized statistics in the Supplemental Matter [29].
Using the second quantized operators above, we introduce the OBC or PBC Hamiltonian
@@
where $n_{i}=\sum_{a}n_{i,a}$ and $n_{i,a}=\psi^{+}_{i,a}\psi^{-}_{i,a}$ are the total and flavor particle densities, respectively. We note that in analogy to Ref. [21], a sum over the internal flavor degrees is performed, making the Hamiltonian flavor-blind. In contrast to Ref. [21], we do include an explicit interaction term which is allowed to be long range for the general discussion below. In the specific examples considered at the end of this letter, the interaction will be limited to nearest neighbors only. For a single mode, we denote the number of possible states with $n$ particles by $d_{n}$ [22].
Hilbert space structureThe first important point is that for flavor-blind bilinear Hamiltonians of the form (2), with operators satisfying the algebra (1), the Hilbert space factorizes.
Factorization theorem: At each lattice site $i$, the local Hilbert space admits a decomposition using the basis states $|n_{i}\rangle$ of the local density operator $n_{i}=\sum_{a}\psi^{+}_{i,a}\psi^{-}_{i,a}$ as
$\mathcal{H}_{i}=\bigoplus_{n=0}^{n_{\rm max}}\mbox{span}\{|n_{i}\rangle\}\otimes\mathcal{F}_{n}$ (3)
where the local flavor space dimension, $\mbox{dim}\,\mathcal{F}_{n}=d_{n}$, is determined by the algebra (1). Because the Hamiltonian (2) sums over the internal flavor index, its action on the internal indices is completely fixed by the quadratic algebra and does not introduce additional flavor dependence; in particular, within each flavor sector, defined by the flavor symmetry algebra commuting with the Hamiltonian, it acts trivially on the corresponding multiplicity space. The full Hilbert space therefore separates into an occupation sector and a flavor sector which is merely a spectator and contributes only degeneracies. A proof of this theorem is given in the End Matter. For OBC, the linear ordering of sites fixes a canonical identification of flavor spaces across different occupation configurations, so the degeneracy is simply the dimension of the flavor space compatible with the occupation number eigenstate. A general state at fixed particle number $N$ is given by $|\Psi\rangle_{N}=\sum_{s}a_{s}|\{n_{i}\}_{s}\rangle\otimes|\phi_{\rm fl}\rangle$ with $\sum_{i}n_{i}=N$ for each $s$. Each occupation number configuration $|\{n_{i}\}\rangle$ has a compatible flavor space $\mathcal{F}(\{n_{i}\})=\bigotimes_{i}\mathcal{F}_{n_{i}}$.
OBC degeneracy rule: The dimension of the flavor space compatible with all occupation configurations entering the superposition $|\Psi\rangle_{N}$ is $D=\mbox{dim}\left[\bigcap_{a_{s}\neq 0}\mathcal{F}(\{n_{i}\}_{s})\right]$.
@@
This all changes when we consider PBC. In this case we can wrap a flavor around the chain thus obtaining cyclic permutations in the order of the flavors. To understand how the eigenstates of the paraparticle Hamiltonian can be constructed from an occupation number Hamiltonian plus flavor degeneracies, one then has to study some basic properties of the cyclic group first. We define $N_{n}$ as the number of particles occupying modes with degeneracy $d_{n}$. Then $N=\sum_{n}N_{n}=\sum_{i}n_{i}$ is the total particle number. We define, furthermore, $M=M(\{N_{n}\})$ as the total number of flavors which actually does get cyclically permuted under PBC. It is important to note that the length of the flavor string $M$ is, in general, not equal to the number of particles $N$. Consider, for example, the case $d_{0}=1$, $d_{1}=2$, $d_{2}=1$, and $d_{n}=0$ for $n>2$. In this case $M=N_{1}$, where $N_{1}$ is the number of singly occupied modes because the vacuum and the doubly occupied modes do not have a flavor index.
@@
with $\delta\equiv\gamma_{q}$. That is, the occupation number wavefunction $|\Psi_{\rm occ}\rangle$ picks up a phase which is equal and opposite to the phase picked up by the flavor part. The eigenspaces of the occupation number Hamiltonian for PBC thus separate into spaces with Peierls phases $\gamma_{q}$ which is equivalent to saying that there is now a flux penetrating the ring. As a consequence, the paraparticle statistics—which is responsible for the phases $\gamma_{q}$—is directly reflected in the eigenspace structure and energies of the occupation number Hamiltonian $H_{\rm occ}$ for PBC.
@@
Example—To demonstrate these general results, we consider from now on example 3 of Ref. [21] with $d_{0}=1$, $d_{1}=m$, and $d_{n}=0$ for $n\geq 2$. In this hardcore case, there is only one type of flavor which is associated with the single occupancy of a mode and every particle carries one of the $m$ possible flavor labels implying $M=M_{1}=N$. A natural realization is the $SU(m)$ Hubbard model with on-site interaction $U\to\infty$ and a nearest-neighbor interaction $V$. For the $SU(2)$ case without $V$ a mapping to free spinless fermions with a twist has already been discussed earlier in the literature [31, 32, 33] but without any reference to parastatistics and in a model-specific manner, whereas Eq. (Parastatistics revealed: Peierls-phase twists and shifted conformal towers in interacting periodic chains) is completely general and relies on the properties of the cyclic group only. In addition to the hopping terms and chemical potential considered in Ref. [21], we include the aforementioned nearest-neighbor density-density interaction with strength $V$. The paraparticle chain can then be mapped onto a spin chain
@@
with $n_{i}=\sum_{\sigma}S^{+}_{i\sigma}S^{-}_{i\sigma}$ where $i$ denotes the sites of the lattice and $\sigma=1,\cdots,m$ is the flavor index. Note that in order to obtain an exact solution, we have chosen the hopping $J$, chemical potential $\mu$, and interaction $V$ to be site independent. The following separation into occupation number and flavor part, however, also holds for site-dependent parameters. For this particular model, the separation into these two parts is achieved easily by embedding the Hilbert space into a larger one, $\mathcal{H}\subset\mathcal{H}^{\rm occ}\otimes\mathcal{H}^{\rm fl}$. Here the local occupation Hilbert space is two-dimensional $H^{\rm occ}_{i}=\mbox{span}(|0\rangle,|1\rangle)$ while the local flavor space is $m$-dimensional, $\mathcal{H}_{i}^{\rm fl}=\mbox{span}(|s_{1}\rangle,\cdots,|s_{m}\rangle)$. To achieve the embedding we simply identify $|0\rangle\equiv|0,s_{1}\rangle$. The Hamiltonian then becomes
@@
for OBC (PBC) with $N$ being the particle number. In a block with $N$ fixed, the chemical potential only contributes a constant but does not affect the eigenvectors. The last term is only present for PBC in which case the Peierls phase is $\gamma_{q}(N)=2\pi q/N$ which can also be distributed uniformly between all bonds. For OBC there is thus no separation into different flux sectors and the dimension of the flavor sector for fixed NN—which determines the degeneracy of each XXZ eigenvalue—is simply $D=m^{N}$. For PBC, we can use the general degeneracy formula (Parastatistics revealed: Peierls-phase twists and shifted conformal towers in interacting periodic chains) with $c_{1}=g=\mbox{gcd}(N,r)$ and $d_{1}=m$ leading to
@@
While the many-body eigenenergies in the interacting case can be obtained by the Bethe ansatz—or by numerical methods for cases where the occupation number Hamiltonian is not integrable—they can be constructed in the non-interacting case, $V=0$, from the single-particle eigenenergies alone. For OBC, these energies are $\varepsilon_{k}=-2J\cos k_{r}-\mu$ with $k_{r}=\frac{\pi r}{L+1}$, and $r=1,\cdots,L$. The many-body energies are then given by $E(\{n_{k}\})=\sum_{k}n_{k}\varepsilon_{k}$ with $n_{k}\in\{0,1\}$, $\sum_{k}n_{k}=N$, $N=1,\cdots,L$ and each energy with $N$ particles has flavor degeneracy $m^{N}$. For PBC, on the other hand, the single-particle eigenvalues are $\varepsilon_{k}(N,q)=-2J\cos k_{r,q}-\mu$ with $k_{r,q}=\frac{2\pi r+\gamma_{q}}{L}$, $r=0,\cdots,L-1$, and $q=0,\cdots,N-1$. The many-body eigenstates are obtained by going through all allowed values of $N,q$ and in each case constructing all possible eigenvalues $E(\{n_{k}\})=\sum_{k}n_{k}\varepsilon_{k}(N,q)$ with $\sum_{k}n_{k}=N$ and with each one of them having an additional degeneracy of $\mbox{dim}\,\mathcal{H}^{\rm fl}_{N,q}$, see Eq. (9).
Bosonization—For the flavor-blind Hamiltonians (2) investigated here, the non-trivial part of the low-energy physics is entirely determined by the occupation number part of the Hamiltonian. This part can often be described by ordinary fermions, making it possible to classify the universal behavior using standard techniques. To be concrete, we continue with our example where the occupation number part of the Hamiltonian with PBC is an XXZ chain with a Peierls phase. For $|V/J|<2$, every sector with $0<N<L$ is gapless and described by a conformal field theory with central charge $c=1$.
@@
Here $e_{\infty}$ is the energy per site in the thermodynamic limit, $L\to\infty$. The first term on the r.h.s. is the well known universal finite-size correction with central charge $c=1$ and $v$ the velocity of the excitations [38, 39]. The second term is the oscillator contribution, with $N_{n}^{R/L}$ counting the occupation of the $n$-th mode for right/left-movers, respectively. The parastatistics enters through the third term. In the low-energy effective theory, the Peierls phase $\gamma_{q}(N)=2\pi q/N$ leads to a persistent current $I(q)=-\frac{\pi vK}{L}(J-q/N)$ with $J=0$ if $q<\lfloor N/2\rfloor$ and $J=1$ if $q>\lfloor N/2\rfloor$. If $q=N/2$, then we have a degeneracy corresponding to equal and opposite persistent currents. In the finite-size spectrum the presence of these persistent currents means that we have conformal towers shifted by $\Delta E=\frac{2\pi vK}{L}\min\left(\frac{q}{N},1-\frac{q}{N}\right)^{2}$ (for $q$ and NqN-q the shift is the same, corresponding to opposite persistent currents) providing a clear signature of the parastatistics in the energy spectrum for PBC. Each of these levels will carry an additional degeneracy of $\mbox{dim}\,\mathcal{H}^{\rm fl}_{N,q}$ due to the flavor sector, see Eq. (9). The parameters $v$ and $K$ (with $K=1$ in the free case) are known exactly from the Bethe ansatz solution of the XXZ model for arbitrary filling.
Thermodynamics—In the thermodynamic limit, the $1/L$-terms due to the boundary conditions become irrelevant. This means that we can start from the open boundary case when considering the $L\to\infty$ limit where the parastatistics only leads to additional degeneracies due to the flavor sector while the occupation number sector is not affected. Returning to our example, this means that the free energy per site is given by ($\beta=1/T$)